# Victor Shea-Jay Huang1,211footnotemark: 1, Le Zhuo1,211footnotemark: 1, Yi Xin2,3, Zhaokai Wang2,4, Peng Gao2, Hongsheng Li122footnotemark: 2 1CUHK, MMLab  2Shanghai AI Laboratory  3NJU  4SJTU

### v1

@@ Abstract
Diffusion Transformers (DiTs) are a powerful yet underexplored class of generative models compared to U-Net-based diffusion models. To bridge this gap, we introduce TIDE (Temporal-aware Sparse Autoencoders for Interpretable Diffusion transformErs), a novel framework that enhances temporal reconstruction within DiT activation layers across denoising steps. TIDE employs Sparse Autoencoders (SAEs) with a sparse bottleneck layer to extract interpretable and hierarchical features, revealing that diffusion models inherently learn hierarchical features at multiple levels (e.g., 3D, semantic, class) during generative pre-training. Our approach achieves state-of-the-art reconstruction performance, with a mean squared error (MSE) of 1e-3 and a cosine similarity of 0.97, demonstrating superior accuracy in capturing activation dynamics along the denoising trajectory. Beyond interpretability, we showcase TIDE’s potential in downstream applications such as sparse activation-guided image editing and style transfer, enabling improved controllability for generative systems. By providing a comprehensive training and evaluation protocol for DiTs, TIDE contributes to the development of more interpretable, transparent, and trustworthy generative models.
@@
Figure 1: TIDE effectively encodes interpretable features across different levels (class, semantic, and 3D) in pre-trained diffusion transformer [7]. This validates that the diffusion model inherently captures and organizes these multi-level concepts by large-scale generative pre-training, enabling it to perform various downstream diffusion tasks effectively (see Sec. 5.4 for details).
1 Introduction
Recent advancements in diffusion models have revolutionized text-to-image generation, with Diffusion Transformers (DiTs) [43] emerging as a scalable alternative to traditional U-Net backbones [48] by leveraging transformer-based architectures. Improvements in sampling efficiency [60] and model scaling [7, 5, 14, 62, 17] .
@@ 1 Introduction
have further elevated their performance to a new level. However, a critical question remains: how do these models function internally? While interpretability techniques have been extensively developed for U-Net frameworks [28, 58], similar efforts for DiTs are absent. More broadly, interpretability research on diffusion models lags significantly behind that of GANs [18, 4, 32], despite the growing prominence of diffusion models in generative modeling. This gap largely stems from diffusion models’ reliance on latent space transformations across various denoising levels, which complicates efforts to analyze their internal mechanisms. In particular, little attention has been paid to understanding latent activations within these models, leaving a significant void in our ability to interpret their processes.
@@
Feature interpretability, which aims to uncover how neural networks encode functionality through internal representations [13], has seen notable success in language models, especially through the use of sparse autoencoders (SAEs) [39]. These tools disentangle polysemantic neurons into sparse, human-interpretable features [6, 37]. However, existing SAE methods face significant limitations when applied to diffusion models, particularly at the level of activation layers. Diffusion models present unique challenges due to their temporal dynamics: unlike language models, which process static tokens in an auto-regressive manner, DiTs iteratively refine latent representations across timesteps, integrating multimodal inputs from text encoders. This iterative nature introduces difficulties such as long token sequences, high token dimensionality, and time-varying activation patterns, especially for high-resolution images. Similarly, although SAEs have proven effective in vision-language models [10] and convolutional networks [19], they encounter distinct challenges in DiTs, including:
@@
To address these issues, we present Temporal-aware SAEs for Interpretable Diffusion transformErs (TIDE), a framework designed to capture temporal variations in activation patterns and extract sparse, interpretable features throughout the denoising process. To ensure high-fidelity feature learning under sparsity constraints, we introduce specialized training strategies, including progressive sparsity scheduling and random sampling augmentation. Additionally, we explore TIDE’s scaling laws and validate its performance using diffusion loss, demonstrating that TIDE achieves state-of-the-art results compared to other SAE-based methods. Moreover, TIDE effectively learns multi-level features (e.g., 3D, semantics, and class), as shown in Fig. 1, highlighting how diffusion models inherently organize such concepts through large-scale generative pre-training. This finding illustrates their ability to be adapted to various downstream tasks, such as depth estimation [20, 25], semantic segmentation [57], and image classification [30, 38]. Additionally, we propose a temporal analysis that isolates and interprets the roles of individual timesteps in the generation process, offering deeper insights into the latent representations of text-conditioned image synthesis. By tailoring SAEs to the dynamics of DiTs, TIDE advances model interpretability, enhances understanding of generative mechanisms, and lays the foundation for more trustworthy and controllable generative systems.
@@
  • TIDE Framework (Sec 4.4): A paradigm that adapts SAEs to DiT via temporal-aware architecture and progressive sparsity scheduling, coupled with training strategies to extract interpretable features from DiT’s latent space while preserving generation quality and balancing sparsity with denoising efficacy.
@@
  • Scaling Law and Evaluation (Sec 5.3): We analyze the scaling law of TIDE, highlighting its ease of training, and compare its downstream loss with other SAEs to demonstrate the superior performance of TIDE.
@@
  • Interpretation (Sec 5.4 5.5): Using TIDE, we demonstrated that diffusion models learn hierarchical features through generative pre-training, enabling efficient transfer to target domains via fine-tuning. Furthermore, through a decoupled analysis of sampling timesteps, we verify and extend existing understanding about DiT’s Coarse-to-Fine generation.
@@
  • Potential Applications (Sec 6): Using PixArt-$\alpha$ [7], we demonstrate the practical utility of TIDE-learned sparse features in tasks like image editing and style transfer.
@@
Figure 2: Overview of our training process: For each Dit, we train its TIDE individually, extracting activation layers from different timesteps each time. The training of the SAE is conducted using both with and without random sampling. For specific details on the loss design, refer to Sec. 4.3.
2 Related work
2.1 Sparse Autoencoders for Interpretability
@@
Mechanistic interpretability [13] aims to understand how neural networks produce outputs by identifying features encoded in neurons [40]. Sparse autoencoders (SAEs)[41] address the polysemantic nature of neurons caused by superposition, where networks encode more features than dimensions[12, 49], by enforcing sparsity, activating only a small subset of neurons per input. SAEs have proven effective in uncovering human-interpretable features [22, 6], decomposing activations in large language models (LLMs)[56, 61], and extracting interpretable concepts in vision-language models such as CLIP[9]. They have also facilitated image editing [10], enabled concept naming [46], and disabled specific generative features [6], demonstrating their potential to mitigate superposition and enhance interpretability across diverse model architectures.
2.2 Interpretability of Diffusion Models
@@
Diffusion models achieve text-to-image generation via iterative denoising, evolving from U-Net architectures [47, 44] to scalable designs like Diffusion Transformer (DiT) [43], with improved sampling efficiency [52, 51, 35]. Interpretability efforts focus on bottleneck layers [28, 42] and cross-attention [55], enabling manipulation of attributes [2] and text encoders [58]. SAEs, applied to CLIP [45, 15] and diffusion models [23, 53, 26], help identify features, reduce hallucinations [24], and analyze activations, linking features to visual semantics.
For more details on related work, please refer to the supplementary material.
3 Preliminaries
3.1 Sparse Autoencoder
@@
3.2 TopK Activation
We employ the $k$-sparse autoencoder [36], which uses the TopK activation function to enforce sparsity. This function selects the $k$ largest latent activations and sets the rest to zero. The function is defined as:
@@ 3.2 TopK Activation
As shown in [16], TopK activation offers several advantages: it directly enforces sparsity by controlling the number of active units, eliminating the need for $\ell_{1}$ regularization; simplifies model comparison by avoiding hyperparameter tuning; improves sparsity-reconstruction trade-offs, especially in larger models; and enhances feature interpretability by zeroing out small activations.
3.3 Diffusion Transformer
Diffusion Transformers (DiT) [43] are transformer-based diffusion models designed for latent space, achieving competitive performance in image generation tasks. Unlike traditional diffusion models that utilize U-Net backbones, DiT employs transformer architectures to process latent patches, offering improved scalability and efficiency. Within the latent diffusion model (LDM) [47] framework, images are first compressed into latent representations via a variational autoencoder (VAE) [27], and then a transformer models the denoising process, reducing computational complexity while preserving output quality.
@@ 3.3 Diffusion Transformer
DiT’s scalable architecture and strong performance on benchmarks like ImageNet position it as a state-of-the-art model for high-resolution image generation.
4 Methodology
4.1 Where to Train Sparse Autoencoders (SAEs)
Recent research on the mechanistic interpretability of diffusion models [2, 3] has revealed that distinct cross-attentaion blocks are responsible for generating particular visual elements, including style and objects. Using these insights, since diffusion models share similar architectures, we selected a representative model, PixArt-XL-2-1024-MS [7] , to extract activation layers for training SAEs. A dedicated SAE was trained for each activation layer. The training process is in Fig. 2
Algorithm 1 Activation Extraction
1:Image $\mathbf{x}_{0}$, Text $\mathbf{p}$, Time step $t$, DiTs $\mathcal{T}$, Target layer $\ell$
@@ 4.1 Where to Train Sparse Autoencoders (SAEs)
4.2 Activations in DiTs
This approach leverages intermediate representations from the generation pipeline for downstream tasks. The activation extraction process is detailed in Algorithm 1.
4.3 Training Loss
The optimization process updates the SAEs’ parameters $\theta$ iteratively using the following formula:
$\theta^{(t+1)}=\theta^{(t)}-\eta_{t}\cdot\text{clip}\left(\mathbf{g}^{(t)},\tau\right),$ (5)
where $\eta_{t}$ denotes the learning rate at step $t$, and $\mathbf{g}^{(t)}$ is the gradient of the training loss with respect to the SAEs’ parameters $\theta$:
$\mathbf{g}^{(t)}=\frac{\partial}{\partial\theta}\left[\ell(x,\hat{x};\theta)+\lambda_{\text{reg}}\cdot R(\theta)\right].$ (6)
The training loss function $\ell(x,\hat{x};\theta)$ consists of multiple components, each capturing a different aspect of the model’s performance. It is defined as:
$\ell(x,\hat{x};\theta)=\mathcal{L}_{\text{mse}}+\mathcal{L}_{\text{cos}}+\mathcal{L}_{\text{temporal}}+\mathcal{L}_{\text{spa}}.$ (7)
Our loss function balances reconstruction,temporal distribution and sparsity.For more details on training, please refer to the supplementary material.
4.4 TIDE: Temporal-Aware SAE
@@
$x_{\text{mod}}=x\cdot(1+\text{scale})+\text{shift},$ (8)
which is initialized from the pre-trained adaptive Layer Normalization in DiT.
@@
5 Experiments
5.1 Training Setup
Train $MSE$ $S_{cos}$ $MSE^{val}$ $S_{cos}^{val}$
SAE $33.3e-$ 3 0.968 $21.23e-$ 2 0.935
SAE(sample) $32.5e-$ 3 0.970 $21.26e-$ 2 0.962
TIDE $33.1e-$ 3 0.970 $21.02e-$ 2 0.944
TIDE(sample) $32.5e-$ 3 0.972 $21.02e-$ 2 0.964
@@ 5.1 Training Setup
Figure 4: (a)(b) Scaling laws of convergence loss with fixed latents n under MSE and cosine similarity loss ($L_{cos}=1-S_{cos}$). (c)(d) Comparison of TIDE and other activation functions: For 73728-d latents, TIDE achieves better trade-offs in diffusion loss against both cosine similarity and MSE.
  • Direct Training: Using 80K prompt-image pairs extracted from CC3M [50] without additional data augmentation.
  • TIDE Training: Incorporating a Temporal-Aware Architecture (Sec 4.4) to enhance the model’s ability to capture temporal dynamics within the activation features.
Evaluation The performance of TIDE was evaluated using cosine similarity, mean squared error (MSE), and sparsity loss (see Sec 4.3 and the supplementary material). These metrics also guided the training process, ensuring that the TIDEs(SAEs) struck a balance between sparsity and reconstruction accuracy.
Hyperparameters The training process employed key hyperparameters (see the supplementary material for details), including token-level sampling rate, cosine annealing learning rate scheduler, and AdamW [34] optimizer with weight decay.
5.2 SAEs / TIDEs Exhibit Superior Reconstruction
We conducted experiments to evaluate the impact of token-level sampling during training and the addition of the temporal-aware architecture in PixArt’s 28 Diffusion Transformers (DiTs) activation layers. The results are summarized in Tab. 1. Our analysis reveals that, while the training strategy without sampling augment did not significantly improve performance on the reconstruction task, it led to a noticeable improvement in cosine similarity on the out-of-distribution (OOD) validation set relative to the training dataset, with only a minor increase in MSE.
This result reveals two key insights: (1) Using the same amount of training data, sampling reduces interference from global image features with only 1/16 of the total tokens, enabling more accurate extraction of critical features and thereby improving both reconstruction precision and generalization. (2) The minimal change in MSE suggests that the data norm remains stable, with only slight directional adjustments. This indicates that the activation distribution is largely preserved, and reconstructed activation layers have minimal impact on subsequent image generation quality.
@@ 5.2 SAEs / TIDEs Exhibit Superior Reconstruction
5.3 Scaling Laws and Downstream Loss of TIDEs
Given the broad capabilities of state-of-the-art (SoTA) diffusion models — distinct from large language models (LLMs) — diffusion models exhibit significantly higher feature dimensionality and process a large number of tokens in each operation. Consequently, we hypothesize that accurately capturing the state of the model requires a large number of sparse features. To address this, we propose two primary approaches for determining the size of the autoencoder and the token budget:
As shown in Fig. 4(a)(b), increasing the number of latent dimensions from $d=4608$ to $7372816d=$ 73728, along with a higher top-k selection, leads to progressively better reconstruction of the activation layers. This improvement is reflected in a MSE reaching the order of $10^{-3}$ and a cosine similarity approaching 0.97. Compared to other diffusion U-Net-based SAEs, as well as most SoTA LLM-based SAEs, the TIDEs trained using our strategy consistently achieve SoTA performance.
We use the downstream diffusion loss to evaluate the reconstruction quality of TIDE. Specifically, the activation layers reconstructed by TIDE are used to replace the original activation layers in DiT, and the diffusion loss is then calculated. As shown in Fig. 4(c)(d),the diffusion loss of TIDE’s reconstructed features is comparable to that of the original model and surpasses other types of SAEs, all while achieving SoTA performance. This highlights TIDE’s effectiveness in preserving features crucial to the generation task.
5.4 Diffusion Really Learned Features
@@
  • 3D-Level: TIDE captures features such as ”far,” ”near,” and ”shadow,” which are indicative of spatial relationships and distinctions between the foreground and background. The activation patterns associated with these features exhibit a high degree of specificity, and training images with similar activations consistently share these characteristics. This consistency underscores the capacity of diffusion models to encode 3D-related concepts critical for downstream tasks like depth estimation.
  • Class-Level: Beyond 3D-level information, TIDE learns class-level representations and their associated sub-features. For example, the class ”Rococo” is characterized by sub-features such as ”house,” ”window,” and ”door,” while ”eagle” is defined by sub-features like ”bird,” ”fly,” and ”beak.” These features demonstrate clear activation patterns localized within distinct regions of the input images, further validating the hierarchical organization of the learned representations.
  • Semantic-Level: Building on the class-level representations, TIDE isolates finer-grained semantic-level features and their corresponding spatial locations within images. For instance, it can precisely identify the ”beak” of an eagle and its position within an image, showcasing the model’s ability to extract localized and semantically rich details.
As shown in Fig.1, these findings confirm that TIDE effectively encodes interpretable features at multiple levels—spanning 3D, class, and semantic representations—within pre-trained DiTs[7]. This analysis underscores the nuanced, hierarchical structure of features learned during large-scale generative pre-training, demonstrating that diffusion models inherently capture and organize multi-level concepts. It provides compelling evidence of their potential to support diverse downstream applications while laying a foundation for future advancements in practical tasks. Consequently, these features enable the models to excel in a variety of downstream tasks. For instance, the ability to capture class-level and semantic-level representations underpins the model’s suitability for tasks like classification [30, 38] and segmentation, as evidenced by zero-shot, unsupervised approaches such as DiffSeg [57]. Similarly, the capability to learn 3D-related concepts empowers diffusion models to tackle depth estimation challenges effectively, as demonstrated by SoTA methods like Marigold [25] for single-image depth prediction and DepthFM [20] for rapid monocular depth estimation.
Figure 5: The generation progress shows that at each timestep, DiT’s activation layers form a directed vector connecting latent features to target features. Abrupt changes in cosine similarity and MSE mark the timestep when contours emerge, signaling the alignment of the latent space with the target manifold. Early modifications aid alignment, while later ones refine details. Using this, we reconstruct latent representations and guide modifications with a directed vector at each step.

### v2

# TIDE : Temporal-Aware Sparse Autoencoders for Interpretable Diffusion Transformers in Image Generation

@@ Abstract
Diffusion Transformers (DiTs) are a powerful yet underexplored class of generative models compared to U-Net-based diffusion architectures. We propose TIDE—Temporal-aware sparse autoencoders for Interpretable Diffusion transformErs—a framework designed to extract sparse, interpretable activation features across timesteps in DiTs. TIDE effectively captures temporally-varying representations and reveals that DiTs naturally learn hierarchical semantics (e.g., 3D structure, object class, and fine-grained concepts) during large-scale pretraining. Experiments show that TIDE enhances interpretability and controllability while maintaining reasonable generation quality, enabling applications such as safe image editing and style transfer.
@@
Figure 1: TIDE effectively encodes interpretable features across different levels (class, semantic, and 3D) in pre-trained diffusion transformer (Chen, Yu et al. 2023). This validates that the diffusion model inherently captures and organizes these multi-level concepts by large-scale generative pre-training, enabling it to perform various downstream diffusion tasks effectively(see Sec.Diffusion Really Learned Features for details).
Introduction
@@
Recent advancements in diffusion models have revolutionized text-to-image generation, with Diffusion Transformers (DiTs) (Peebles and Xie 2022) emerging as a scalable alternative to traditional U-Net backbones (Ronneberger, Fischer, and Brox 2015) by leveraging transformer-based architectures. Improvements in sampling efficiency (Yi et al. 2024) and model scaling (Chen, Yu et al. 2023; Esser, Kulal et al. 2024; Zhuo, Du et al. 2024; Gao, Zhuo et al. 2024) have further elevated their performance to a new level. However, a critical question remains: how do these models function internally? While interpretability techniques have been extensively developed for U-Net frameworks (Kwon, Jeong, and Uh 2023; Toker et al. 2024), similar efforts for DiTs are absent. This is primarily due to the fact that most existing methods are tailored to convolutional architectures and depend on spatially structured feature maps, which are not directly available in the token-based representation of DiTs. More broadly, interpretability research on diffusion models lags significantly behind that of GANs (Goodfellow et al. 2014; Bau et al. 2018; Ling et al. 2021), despite the growing prominence of diffusion models in generative modeling. This gap largely stems from diffusion models’ reliance on latent space transformations across various denoising levels, which complicates efforts to analyze their internal mechanisms. In particular, little attention has been paid to understanding latent activations within these models, leaving a significant void in our ability to interpret their processes.
@@
Feature interpretability, which aims to uncover how neural networks encode functionality through internal representations (Elhage, Nanda et al. 2021), has seen notable success in language models, especially through the use of sparse autoencoders (SAEs) (Ng et al. 2011). These tools disentangle polysemantic neurons into sparse, human-interpretable features (Bricken et al. 2023; Marks et al. 2024). However, existing SAE methods face significant limitations when applied to diffusion models, particularly at the level of activation layers. Diffusion models present unique challenges due to their temporal dynamics: unlike language models, which process static tokens in an auto-regressive manner, DiTs iteratively refine latent representations across timesteps, integrating multimodal inputs from text encoders. This iterative nature introduces difficulties such as long token sequences, high token dimensionality, and time-varying activation patterns, especially for high-resolution images. Similarly, although SAEs have proven effective in vision-language models (Daujotas 2024a) and convolutional networks (Gorton 2024), they encounter distinct challenges in DiTs, including:
@@
To address these issues, we present Temporal-aware SAEs for Interpretable Diffusion transformErs (TIDE), a framework designed to capture temporal variations in activation patterns and extract sparse, interpretable features throughout the denoising process. To ensure high-fidelity feature learning under sparsity constraints, we introduce specialized training strategies, including progressive sparsity scheduling and random sampling augmentation. Additionally, we explore TIDE’s scaling laws and validate its performance using diffusion loss, demonstrating that TIDE achieves advanced results compared to other SAE-based methods. Moreover, TIDE effectively learns multi-level features (e.g., 3D, semantics, and class), as shown in Fig. 1, highlighting how diffusion models inherently organize such concepts through large-scale generative pre-training. This finding illustrates their ability to be adapted to various downstream tasks, such as depth estimation (Gui et al. 2024; Ke et al. 2024), semantic segmentation (Tian et al. 2024), and image classification (Li et al. 2023; Mukhopadhyay et al. 2023). Additionally, we propose a temporal analysis that isolates and interprets the roles of individual timesteps in the generation process, offering deeper insights into the latent representations of text-conditioned image synthesis. By tailoring SAEs to the dynamics of DiTs, TIDE advances model interpretability, enhances understanding of generative mechanisms, and lays the foundation for more trustworthy and controllable generative systems.
@@
  • TIDE Framework: A paradigm that adapts SAEs to DiT via temporal-aware architecture and progressive sparsity scheduling, coupled with training strategies to extract interpretable features from DiT’s latent space while preserving generation quality and balancing sparsity with denoising efficacy.
@@
  • Scaling Law and Evaluation: We analyze the scaling law of TIDE, highlighting its ease of training, and compare its downstream loss with other SAEs to demonstrate the superior performance of TIDE.
@@
  • Interpretation: Using TIDE, we demonstrated that diffusion models learn hierarchical features through generative pre-training, enabling efficient transfer to target domains via fine-tuning. Furthermore, through a decoupled analysis of sampling timesteps, we verify and extend existing understanding about DiT’s Coarse-to-Fine generation.
@@
  • Potential Applications: Using PixArt-$\alpha$ (Chen, Yu et al. 2023) and other diffusion models, we demonstrate the practical utility of TIDE-learned sparse features in tasks like safe image editing and style transfer.
@@
Figure 2: Overview of our training process: For each Dit, we train its TIDE (Embed SAE within the TA (temporal-aware) architecture) individually, extracting activation layers from different timesteps each time. The training of the SAE is conducted using both with and without random sampling. For specific details on the loss design, refer to supporting material.
Related work
Sparse Autoencoders for Interpretability
@@
Mechanistic interpretability (Elhage, Nanda et al. 2021) aims to understand how neural networks produce outputs by identifying features encoded in neurons (Olah et al. 2020). Sparse autoencoders (SAEs)(Olshausen and Field 1997) address the polysemantic nature of neurons caused by superposition, where networks encode more features than dimensions(Elhage et al. 2022; Sharkey, Braun, and beren 2022), by enforcing sparsity, activating only a small subset of neurons per input. SAEs have proven effective in uncovering human-interpretable features (Huben et al. 2024; Bricken et al. 2023), decomposing activations in large language models (LLMs)(Templeton 2024; Yun et al. 2021), and extracting interpretable concepts in vision-language models such as CLIP(Daujotas 2024b). They have also facilitated image editing (Daujotas 2024a), enabled concept naming (Rao et al. 2024), and disabled specific generative features (Bricken et al. 2023), demonstrating their potential to mitigate superposition and enhance interpretability across diverse model architectures.
Interpretability of Diffusion Models
@@
Diffusion models achieve text-to-image generation via iterative denoising, evolving from U-Net architectures (Rombach et al. 2022a; Podell et al. 2023) to scalable designs like Diffusion Transformer (DiT) (Peebles and Xie 2022), with improved sampling efficiency (Song and Ermon 2019; Song, Meng, and Ermon 2020; Lu et al. 2022). Interpretability efforts focus on bottleneck layers (Kwon, Jeong, and Uh 2023; Park et al. 2023) and cross-attention (Tang et al. 2022), enabling manipulation of attributes (Basu et al. 2023) and text encoders (Toker et al. 2024). SAEs, applied to CLIP (Radford et al. 2021; Fry 2024) and diffusion models (Ijishakin et al. 2024; Surkov et al. 2024; Kim, Thomas, and Ghadiyaram 2024; Shabalin et al. 2025), help identify features, reduce hallucinations (Jiang et al. 2024), and analyze activations, linking features to visual semantics. For more details on related work, please refer to the supplementary material.
Preliminaries
Sparse Autoencoder
@@
TopK Activation
We employ the $k$-sparse autoencoder (Makhzani and Frey 2013), which uses the TopK activation function to enforce sparsity. This function selects the $k$ largest latent activations and sets the rest to zero. The function is defined as:
@@
As shown in (Gao et al. 2024), TopK activation offers several advantages: it directly enforces sparsity by controlling the number of active units, eliminating the need for $\ell_{1}$ regularization; simplifies model comparison by avoiding hyperparameter tuning; improves sparsity-reconstruction trade-offs, especially in larger models; and enhances feature interpretability by zeroing out small activations.
Diffusion Transformer
Diffusion Transformers (DiT)(Peebles and Xie 2022) are transformer-based diffusion models operating in latent space, offering competitive image generation performance. Unlike traditional U-Net-based models, DiT leverages transformers to process latent patches, enhancing scalability and efficiency. In the latent diffusion framework (LDM)(Rombach et al. 2022a), images are first encoded into latents via a variational autoencoder (VAE) (Kingma, Welling et al. 2013), and then a transformer models the denoising process, reducing computation while maintaining quality.
@@
DiT’s scalable architecture and strong performance on benchmarks like ImageNet position it as a advanced model for high-resolution image generation.
Methodology
Where to Train Sparse Autoencoders (SAEs)
Recent research on the mechanistic interpretability of diffusion models (Basu et al. 2023, 2024) has revealed that distinct cross-attentaion blocks are responsible for generating particular visual elements, including style and objects. Using these insights, since diffusion models share similar architectures, we selected a representative model, PixArt-XL-2-1024-MS (Chen, Yu et al. 2023) , to extract activation layers for training SAEs. A dedicated SAE was trained for each activation layer. The training process is in Fig. 2
Activations in DiTs
This approach leverages intermediate representations from the generation pipeline for downstream tasks. The activation extraction process is detailed in algorithm below:
Algorithm 1 Extract Activations from DiTs
1:Image $\mathbf{x}_{0}$, Text Prompt $\mathbf{p}$, Time step $t$, DiTs $\mathcal{T}$, Target layer $\ell$
@@
TIDE: Temporal-Aware SAE
@@
$x_{\text{mod}}=x\cdot(1+\text{scale}(t))+\text{shift}(t),$ (5)
where both scale(t) and shift(t) are functions of the current timestep $t$, modeled via a lightweight MLP initialized from the pre-trained adaptive Layer Normalization in DiT. This modulation allows the model to adaptively align activations across different diffusion timesteps.
@@
Experiments
Training Setup
  • Direct Training: Using 80K prompt-image pairs extracted from CC3M (Sharma et al. 2018) without augmentation.
  • TIDE Training: Incorporating a Temporal-Aware Architecture to enhance the model’s ability to capture temporal dynamics within the activation features.
Evaluation The performance of TIDE was evaluated using cosine similarity, mean squared error (MSE), and sparsity loss (see the supplementary material). These metrics also guided the training process, ensuring that the TIDEs(SAEs) struck a balance between sparsity and reconstruction accuracy. A validation set of 500 prompt-image pairs is sampled from the CC3M dataset and is not used during training. For TIDE (SAE) models trained solely on the training data, this validation set is considered out-of-distribution (OOD).
SAEs / TIDEs Exhibit Superior Reconstruction
@@
Train MSEMSE $S_{cos}$ $MSE^{val}$ $S_{cos}^{val}$
SAE 3.3e33.3e-3 0.968 1.23e21.23e-2 0.935
SAE(sample) 2.5e32.5e-3 0.970 1.26e21.26e-2 0.962
TIDE 3.1e33.1e-3 0.970 1.02e21.02e-2 0.944
TIDE(sample) 2.5e32.5e-3 0.972 1.02e21.02e-2 0.964
We conducted experiments to evaluate the impact of token-level sampling during training and the integration of temporal-aware architecture into PixArt’s 28 Diffusion Transformer (DiT) activation layers. As shown in Tab. 1, selecting only 1/16 of the total tokens during training significantly improves reconstruction accuracy on the training set (columns 1–2), while introducing only a minor increase in MSE. Notably, this strategy also leads to a marked improvement in cosine similarity on the out-of-distribution (OOD) validation set (column 4), suggesting enhanced generalization. In contrast, training without token sampling shows limited gains in reconstruction performance but fails to improve OOD alignment.
This result reveals two key insights: (1) Token-level sampling improves learning efficiency by focusing on more informative local features, while also mitigating the impact of global noise—leading to better generalization. (2) The negligible change in MSE indicates that the norm of the activation distribution remains stable, implying that the reconstruction process preserves essential structural information and does not adversely affect downstream image generation quality.
@@
Scaling Laws and Downstream Loss of TIDEs
Figure 4: (a)(b) Scaling laws of convergence loss with fixed latents n under MSE and cosine similarity loss ($L_{cos}=1-S_{cos}$). (c)(d) Comparison of TIDE and other activation functions: For 73728-d latents, TIDE achieves better trade-offs in diffusion loss against both cosine similarity and MSE.Based on these experiments, we selected top-k values in the range of 1024 to 4096 and set 16d = 73,728 as the parameters for subsequent image editing tasks.
Given the broad capabilities of advanced diffusion models — distinct from large language models (LLMs) — diffusion models exhibit significantly higher feature dimensionality and process a large number of tokens in each operation. Consequently, we hypothesize that accurately capturing the state of the model requires a large number of sparse features. To address this, we propose two primary approaches for determining the size of the autoencoder and the token budget:
As shown in Fig. 4(a)(b), increasing the number of latent dimensions from $d=4608$ to $16d=73728$, along with a higher top-k selection, leads to progressively better reconstruction of the activation layers. This improvement is reflected in a MSE reaching the order of $10^{-3}$ and a cosine similarity approaching 0.97. Compared to other diffusion U-Net-based SAEs, as well as most advanced LLM-based SAEs, the TIDEs trained using our strategy consistently achieve advanced performance.
We use the downstream diffusion loss to evaluate the reconstruction quality of TIDE. Specifically, the activation layers reconstructed by TIDE are used to replace the original activation layers in DiT, and the diffusion loss is then calculated. As shown in Fig. 4(c)(d),the diffusion loss of TIDE’s reconstructed features is comparable to that of the original model and surpasses other types of SAEs, all while achieving advanced performance. This highlights TIDE’s effectiveness in preserving features crucial to the generation task.
Diffusion Really Learned Features